# Measurement-induced phase transitions for free fermions in a quasiperiodic potential

### v1

@@ I Introduction
It is worth noting that the MIPT in free-particle systems is one of the most actively investigated subjects [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. Recently, the absence of MIPTs in free fermions has been reported in one-dimensional (1D) tight-binding models on the spatially homogeneous systems by analyzing the nonlinear sigma models as an effective field theory [68]. On the other hand, the existence of MIPTs has been clarified in free fermions above one-dimension [69, 70, 71], in free fermions and bosons with long-range couplings [72, 73, 74], and in disordered free fermions [75, 76]. Therefore, many open questions remain about the universality and mechanisms underlying MIPTs in free fermions with yet unnoticed structures.
As most of the existing studies on MIPTs have considered homogeneous and disordered systems, it is timely to focus on another system with incommensurate structures, or quasiperiodic systems, which are characterized by aperiodic but long-range ordered structures. In quantum circuits, several studies have analyzed such problems, where the measurement profiles are modulated to follow the quasiperiodicity with respect to sites [38], strengths [39], and spatiotemporal distributions [14, 31]. Thus, quasiperiodic systems under random measurement offer a rich possibility for studying MIPTs.
However, dynamics under continuous measurement, where quasiperiodicity is imposed on parameters in the system Hamiltonian such as potentials and lattice configurations, remains largely unexplored. This setup is one of the most fundamental platforms in condensed matter physics, and it is now possible to control quasiperiodic structures in optical lattices by using ultracold atomic gases, such as in 1D [77, 78, 79, 80, 81, 82, 83], pentagonal [84, 85, 86], and octagonal [87, 88, 89] systems. Experimentally, these structures are created by superimposing several standing waves with incommensurate wavelengths or angles. Importantly, quasiperiodic systems exhibit unique localization phenomena distinct from both homogeneous and disordered systems. For example, the Aubry-André-Harper (AAH) model [90, 91], which describes free fermions in a 1D lattice under a quasiperiodic potential, displays transitions in all eigenstate wave functions between extended and localized states with increasing the potential strength [92]. These localization properties also manifest in the scaling behavior of steady-state entanglement entropy as the volume law and area law, respectively [93, 94]. Therefore, we ask for the following question: Are there MIPTs in continuously monitored free fermions under quasiperiodic potentials and how does the measurement affect the entanglement scaling?
@@ II Model and method
where $\rho_{A}$ is the reduced density matrix of subsystem $A$. We set the subsystem to ${A}=[1,L/2]$, where $L$ is the system size. In our simulations, we set an initial state to the Néel state, whose entanglement entropy is zero. When time evolves, $S$ increases, and eventually saturates in the long-time limit. (For the transient behavior before reaching the steady-state, see Appendix C).
@@ IV.1 Phase diagram and steady-state entanglement entropy
Figure 4: System-size dependence of the entanglement entropy for (a) $V/J=1$, (b) $2$, (c) $2.2$, and (d) $3$. The horizontal axis is set to be a logarithmic scale. The color map represents the values of the measurement strength $\gamma$. The entanglement entropy at $\gamma/J=0$ is depicted by the dashed line. In this simulation, we use $(N_{\rm pot},N_{\rm traj})=(1,100)$.
Figure 5: Finite-size scaling analysis of the entanglement entropy in the systems with $L=150,200,\cdots,500$ when (a) $V/J=0.1$, (b) 0.2, (c) 0.5, (d) 1, and (e) 1.5. In this simulation, we use $(N_{\rm pot},N_{\rm traj})=(1,100)$.
@@ IV.2 Finite-size scaling analysis
We use the finite-size scaling analysis for the cases with several $V/J$. In Fig. 5, we find that the appropriate values of $\gamma_{c}$ give fairly good scaling plots. This is consistent with the scaling formula (5). To determine the transition point and its error bar, we minimize the cost function [75], where the error bar of $\gamma_{c}$ is estimated from the range within twice of its minimum value. The obtained results are shown as the circles with error bars in Fig. 3, and summarized in Table 1. We find that the critical measurement strength $\gamma_{c}$ decreases monotonically as $V$ increases. This result is different from the non-monotonic phase boundary reported in disordered free fermions [75]. This may lead to measurement-induced properties unique to quasiperiodic systems that are not seen in monitored dynamics under uniform disorder.
@@ V.1 Connected correlation functions
Figure 6: Connected correlation function $\overline{C}(r)$ for the potential strength (a) $V/J=0.5$, (b) 2, and (c) 2.2. The curves with light, medium, and dark colors represent the results for system sizes $L=100$, 300, and 500, respectively. Gray lines show the decay $\propto[(L/\pi)\sin(\pi r/L)]^{-2}$, which approaches $r^{-2}$ in the thermodynamic limit $L\rightarrow\infty$. In this simulation, we use $(N_{\rm pot},N_{\rm traj})=(46,23)$.
@@ V.2 Wave functions
We examine the localization properties of single-particle wave functions. We recall that, in the unitary limit $\gamma/J=0$, the eigenstate wave function is extended (localized) when $V/J<2$ ($V/J>2$), and critical at $V/J=2$. Accordingly, distinct single-particle properties in the long-time limit have been observed in Fig. 1. Figure 8 shows the single-partilce wave function under the continuous monitoring, where we shift the amplitudes in the spatial direction such that their maximum values are positioned at $j=0$. The amplitudes are then averaged over multiple trajectories, resulting in the clear behavior of the localized wave functions. When $V/J=1$, the single-particle wave function is similar to that observed in the unitary limit since a finite amplitude exists at each site. This may imply that this quantity is not appropriate for distinguishing the volume-law phase in the unitary limit and the logarithmic-law phase under measurement. However, a drastic change appears for $2\leq V/J\leq 5$. In particular, the wave function away from the center of the system is changed by the continuous measurement; its amplitude is zero when $\gamma/J=0$, while it is finite when $\gamma/J=0.01$. Therefore, the single-particle wave function may capture the measurement-induced logarithmic-law phase. When $V/J=10$, the localization of the wave function becomes more pronounced, as shown in the inset of Fig. 8. This indicates that the system is in the area-law phase II characterized by the localized wave functions. Although determining the phase transition point remains challenging, this analysis provides insight into these phases.
@@ VI Summary
We have studied the dynamics under continuous measurement for free fermions in a quasiperiodic potential by using the AAH model. We have revealed the existence of the robust logarithmic phase for weak potential strength. It has been found that a further increase in the measurement strength induces the MIPT to the area-law phase. For strong potential strength, we have demonstrated that MIPTs vanish and the area-law phase is stable irrespective of the measurement strength. We have further analyzed connected correlation functions and single-particle wave functions to support the above results: connected correlation functions exhibit an algebraic decay and amplitudes of single-particle wave functions tend to a constant at the edges in the logarithmic-law phase. These findings highlight the distinct behavior of MIPTs in quasiperiodic systems compared to the unitary limit and lead to the understanding of entanglement dynamics in quasiperiodic quantum systems.
@@ Acknowledgements.
Parts of the numerical calculations are performed in the supercomputing systems in ISSP, the University of Tokyo. T.M. was supported by JST SPRING, Japan Grant Numbers JPMJSP2106 and JPMJSP2180. This work was supported by Grant-in-Aid for Scientific Research from JSPS, KAKENHI Grant Numbers JP23K19031 (K.Y.) and JP22K03525 (A.K.). K.Y. was also supported by Toyota RIKEN Scholar Program, Murata Science and Education Foundation, Public Promoting Association Kura Foundation, Hirose Foundation, the Precise Measurement Technology Promotion Foundation, and the Fujikura Foundation.

### v2

@@ I Introduction
It is worth noting that the MIPT in free-particle systems is one of the most actively investigated subjects [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]. Recently, the absence of MIPTs in free fermions has been reported in one-dimensional (1D) tight-binding models on the spatially homogeneous systems by analyzing the nonlinear sigma models as an effective field theory [72, 73, 74]. On the other hand, the existence of MIPTs has been clarified in free fermions above one-dimension [75, 76, 77], in free fermions and bosons with long-range couplings [78, 79, 80], and in disordered free fermions [81, 82]. Therefore, many open questions remain about the universality and mechanisms underlying MIPTs in free fermions with yet unnoticed structures.
As most of the existing studies on MIPTs have considered homogeneous and disordered systems, it is timely to focus on another system with incommensurate structures, or quasiperiodic systems, which are characterized by aperiodic but long-range ordered structures. In quantum circuits, several studies have analyzed such problems, where the measurement profiles are modulated to follow the quasiperiodicity with respect to sites [38], strengths [39], and spatiotemporal distributions [14, 31]. In the context of non-Hermitian quasiperiodic systems, several properties of entanglement entropy have been also studied [83, 84]. Thus, quasiperiodic systems under random measurement offer a rich possibility for studying MIPTs.
However, dynamics under continuous measurement, where quasiperiodicity is imposed on parameters in the system Hamiltonian such as potentials and lattice configurations, remains largely unexplored. This setup is one of the most fundamental platforms in condensed matter physics, and it is now possible to control quasiperiodic structures in optical lattices by using ultracold atomic gases, such as in 1D [85, 86, 87, 88, 89, 90, 91, 92, 93], pentagonal [94, 95, 96, 97], and octagonal [98, 99, 100, 101, 102] systems. Experimentally, these structures are created by superimposing several standing waves with incommensurate wavelengths or angles. Importantly, quasiperiodic systems exhibit unique localization phenomena distinct from both homogeneous and disordered systems. For example, the Aubry-André-Harper (AAH) model [103, 104], which describes free fermions in a 1D lattice under a quasiperiodic potential, displays transitions in all eigenstate wave functions between extended and localized states with increasing the potential strength [105]. These localization properties also manifest in the scaling behavior of steady-state entanglement entropy as the volume law and area law, respectively [106, 107]. Therefore, we ask for the following question: Are there MIPTs in continuously monitored free fermions under quasiperiodic potentials and how does the measurement affect the entanglement scaling?
@@ II Model and method
where $\rho_{A}$ is the reduced density matrix of subsystem $A$. We set the subsystem to A=[1,L/2]{A}=[1,L/2], where $L$ is the system size. In our simulations, we set an initial state to the Néel state, whose entanglement entropy is zero. When time evolves, $S$ increases, and eventually saturates in the long-time limit. (For the transient behavior before reaching the steady-state, see Appendix C).
@@ IV.1 Phase diagram and steady-state entanglement entropy
Figure 4: System-size dependence of the entanglement entropy for (a) $V/J=1$, (b) 22, (c) $2.2$, and (d) 33. The horizontal axis is set to be a logarithmic scale. The color map represents the values of the measurement strength $\gamma$. The entanglement entropy at $\gamma/J=0$ is depicted by the dashed line. In this simulation, we use $(N_{\rm pot},N_{\rm traj})=(1,100)$.
Figure 5: Finite-size scaling analysis of the entanglement entropy in the systems with $L=150,200,\cdots,500$ when (a) $V/J=0.1$, (b) 0.2, (c) 0.5, (d) 1, (e) 1.5, and (f) 2. In this simulation, we use $(N_{\rm pot},N_{\rm traj})=(1,100)$.
@@ IV.2 Finite-size scaling analysis
We use the finite-size scaling analysis for the cases with several $V/J$. In Fig. 5, we find that the appropriate values of $\gamma_{c}$ give fairly good scaling plots. This is consistent with the scaling formula (5). To determine the transition point and its error bar, we minimize the cost function [81], where the error bar of $\gamma_{c}$ is estimated from the range within twice of its minimum value. The obtained results are shown as the circles with error bars in Fig. 3, and summarized in Table 1. We find that the critical measurement strength $\gamma_{c}$ decreases monotonically as $V$ increases. This result is different from the non-monotonic phase boundary reported in disordered free fermions [81]. This may lead to measurement-induced properties unique to quasiperiodic systems that are not seen in monitored dynamics under uniform disorder. We have also tested the finite-size scaling with $\left.\overline{S}\right|_{\gamma}-\left.\overline{S}\right|_{\gamma_{c}}=F[(\gamma-\gamma_{c})(\log L)]$, and confirmed that the data do not change the results qualitatively, although it may affect the quantitative difference of the transition points.
@@ V.1 Connected correlation functions
Figure 6: Connected correlation function $\overline{C}(r)$ for the potential strength (a) $V/J=0.5$, (b) 2, and (c) $2.$ The curves with light, medium, and dark colors represent the results for system sizes $L=100$, 300, and 500, respectively. Gray lines show the decay $\propto[(L/\pi)\sin(\pi r/L)]^{-2}$, which approaches $r^{-2}$ in the thermodynamic limit $L\rightarrow\infty$. In this simulation, we use $(N_{\rm pot},N_{\rm traj})=(46,23)$.
@@ V.2 Wave functions
We examine the localization properties of single-particle wave functions. We recall that, in the unitary limit $\gamma/J=0$, the eigenstate wave function is extended (localized) when $V/J<2$ ($V/J>2$), and critical at $V/J=2$. Accordingly, distinct single-particle properties in the long-time limit have been observed in Fig. 1. Figure 8 shows the single-particle wave function under the continuous monitoring, where we shift the amplitudes in the spatial direction such that their maximum values are positioned at $j=0$. The amplitudes are then averaged over multiple trajectories, resulting in the clear behavior of the localized wave functions. When $V/J=1$, the single-particle wave function is similar to that observed in the unitary limit since a finite amplitude exists at each site. This may imply that this quantity is not appropriate for distinguishing the volume-law phase in the unitary limit and the logarithmic-law phase under measurement. However, a drastic change appears for $2\leq V/J\leq 5$. In particular, the wave function away from the center of the system is changed by the continuous measurement; its amplitude is zero when $\gamma/J=0$, while it is finite when $\gamma/J=0.01$. Therefore, the single-particle wave function may capture the measurement-induced logarithmic-law phase. When $V/J=10$, the localization of the wave function becomes more pronounced, as shown in the inset of Fig. 8. This indicates that the system is in the area-law phase II characterized by the localized wave functions. Furthermore, as given in Appendix D, the inverse participation ratio (IPR) quantitatively supports the validity of this discussion. Although determining the phase transition point remains challenging, this analysis provides insight into these phases.
@@ VI Summary
We have studied the dynamics under continuous measurement for free fermions in a quasiperiodic potential by using the AAH model. We have revealed the existence of the robust logarithmic phase for weak potential strength. It has been found that a further increase in the measurement strength induces the MIPT to the area-law phase. For strong potential strength, we have demonstrated that MIPTs vanish and the area-law phase is stable irrespective of the measurement strength. We have further analyzed connected correlation functions and single-particle wave functions to support the above results: connected correlation functions exhibit an algebraic decay and amplitudes of single-particle wave functions tend to a constant at the edges in the logarithmic-law phase. These findings highlight the distinct behavior of MIPTs in quasiperiodic systems compared to the unitary limit and lead to the understanding of entanglement dynamics in quasiperiodic quantum systems. We note that the destruction of localized or insulating states under dephasing has been studied in the context of dephasing-enhanced transport [117, 118, 119, 120], which has been also observed in quasiperiodic systems [121], where weak (strong) dephasing delocalizes (localizes) the wave function when the potential strength is large enough. This is similar to the behavior of the wave function for $2<V/J<2.3$ in the AAH model under continuous measurement, and investigating the effect of quasiperiodicity on the dynamics described by Lindblad equations is also intresting.
@@ Acknowledgements.
Parts of the numerical calculations are performed in the supercomputing systems in ISSP, the University of Tokyo. T.M. was supported by JST SPRING, Japan Grant Numbers JPMJSP2106 and JPMJSP2180. This work was supported by Grant-in-Aid for Scientific Research from JSPS, KAKENHI Grant Numbers JP23K19031, JP25K17327 (K.Y.), and JP22K03525 (A.K.). K.Y. was also supported by Murata Science and Education Foundation, Hirose Foundation, the Precise Measurement Technology Promotion Foundation, and the Fujikura Foundation.
Data availability
The data that support the findings of this article are openly available [124].
@@ Appendix C Dynamics of the entanglement entropy under continuous measurement
Appendix D Inverse participation ratio
To quantitatively characterize the localization of the wave function, we employ IPR as one of the standard measures. The IPR is defined as
[…]
Figure 11: The IPR of the single-particle wave function. The calculation is performed for $L=100$, where the minimum value of IPR is $I=0.01$.
We now calculate IPR to quantitatively evaluate the localization strength of the wave functions shown in Figs. 1 and 8. The results are presented in Fig. 11. In the unitary limit ($\gamma/J=0$), IPR increases sharply when the potential strength exceeds $V/J=2$. This result is consistent with the fact that the wave function becomes localized above $V/J=2$. For $2\leq V/J\leq 5$ with the measurement strength $\gamma/J=0.01$, we see that IPR is suppressed compared to the unitary limit. This result provides quantitative support for the fact that the wave function away from the center is finite under measurement, while it becomes zero in the unitary limit. Furthermore, we find that IPR is enhanced for $\gamma/J=0.01$ as the potential strength is increased. This is consistent with our discussion that the wave function becomes more localized for strong potential strength such as $V/J=10$. These results quantitatively characterize the localization behavior shown in Sec. V.2.